Calculating Unemployment Given the following information about a country’s population, answer the questions which follow. Make...

  1. Calculating Unemployment

Given the following information about a country’s population, answer the questions which follow. Make sure you show all of your calculations.

1987 1992

Non-institutionalized, civilian,           182,753,000                191,576,000

Adult population

Employed persons                              112,440,000                117,598,000

Unemployed persons                          7,425,000                    9,384,000

  1. What was the size of the civilian labor force in 1987? ____________
  2. What was the size of the civilian labor force in 1992? ____________
  3. What was the unemployment rate in 1987? ___________
  4. What was the unemployment rate in 1992? ___________
  5. What was the employment rate in 1987? __________
Answer #1

a. Labor force in 1987 = Employed persons + Unemployed persons = 112,440,000 + 7,425,000 = 119,865,000

b. Labor force in 1992 = Employed persons + Unemployed persons = 117,598,000 + 9,384,000 = 126,982,000

c. Unemployment rate in 1987 = (Unemployed persons/Labor force in 1987)*100 = (7,425,000/119,865,000)*100 = 6.19%

d. Unemployment rate in 1992 = (Unemployed persons/Labor force in 1992)*100 = (9,384,000/126,982,000)*100 = 7.39%

e. Employment rate in 1987 = (Employed persons/Labor force in 1987)*100 = (112,440,000/119,865,000)*100 = 93.81%